OpenStudy (anonymous):

[4_06]Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (y1066380):

y = -3x + 1

OpenStudy (y1066380):

Just put (2,-1) into all the equations and try

OpenStudy (anonymous):

D. y+5 =-3(x-2) y = -3x+6-5 y = -3x + 1 :))

OpenStudy (anonymous):

[4_07]Choose the equation of the line passing through the point (4, -2) and perpendicular to y = one thirdx - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (y1066380):

y = -3x + 10

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

A RIGHT?[4_01]Which of the following is the slope between the points (-1, -5) and (7, -3)? 4 one fourth -one fourth -4

OpenStudy (y1066380):

one fourth

OpenStudy (y1066380):

Slope = [-5-(-3)]/(-1-7) =1/4

OpenStudy (anonymous):

[4_06]Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1)

OpenStudy (y1066380):

C

OpenStudy (y1066380):

all correct?

OpenStudy (anonymous):

A. Vertical lines have undefined slopes..

OpenStudy (anonymous):

[4_03]Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9. and last but not least. xD

OpenStudy (anonymous):

y = -6x+9 :))

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

anytime.. :))

OpenStudy (anonymous):

what course are you taking? just curious.. :)